# Counting with the quantum alternating operator ansatz

### v1

@@ Abstract
We introduce a variational algorithm based on the quantum alternating operator ansatz (QAOA) for the approximate solution of computationally hard counting problems. Our algorithm, dubbed VQCount, is based on the equivalence between random sampling and approximate counting and employs QAOA as a solution sampler. We first prove that VQCount improves upon previous work by reducing exponentially the number of samples needed to obtain an approximation within a multiplicative factor of the exact count. Using tensor network simulations, we then study the typical performance of VQCount with shallow circuits on synthetic instances of two #P-hard problems, positive #NAE3SAT and positive #1-in-3SAT. We employ the original quantum approximate optimization algorithm version of QAOA, as well as the Grover-mixer variant which guarantees a uniform solution probability distribution. We observe a tradeoff between QAOA success probability and sampling uniformity, which we exploit to achieve an exponential gain in efficiency over naive rejection sampling. Our results highlight the potential and limitations of variational algorithms for approximate counting.
1 Introduction
@@
These considerations motivate us to turn our attention to counting problems. This class of problems asks how many solutions there are for a given instance of a decision problem. Counting problems lie higher in the computational complexity hierarchy than the optimization problems commonly targeted with near-term quantum algorithms. Counting problems arise across many disciplines, with applications in probabilistic reasoning [6, 7, 8], network reliability [9, 10], statistical physics [11], and artificial intelligence [12]. For decision problems whose solutions are verifiable in time that scales polynomially with the size of the problem instance, that is, in polynomial time, the corresponding task of counting solutions belongs to complexity class #P. The class #P, like NP, contains a subset of hard problems which are unlikely to admit exact polynomial-time solutions. Naturally, the counting counterpart of every NP-hard problem in NP is #P-hard. In the quantum domain, counting was originally formulated as an application of amplitude estimation [13, 14, 15], which uses the Grover iteration [16] as a subroutine. More recently, VQAs based on QAOA for (weighted) counting [17] and enumeration [18] were introduced. These algorithms are parameterized by the success rate of QAOA circuits, i.e., the probability of a measurement yielding a solution.
The improbability of exact polynomial-time algorithms for #P-hard problems has motivated the development of approximate counting algorithms, which sacrifice exactness for performance. An example is the celebrated algorithm by Jerrum, Valiant, and Vazirani (JVV) [19], which establishes an equivalence between approximate counting and random sampling of solutions. The JVV algorithm approximately solves #P-hard problems that are self-reducible if solutions can be sampled sufficiently uniformly (the term “self-reducible” will be defined in Sec. 3).
In this work, we introduce an algorithm for approximate counting with VQAs. Concretely, we show how the quantum alternating operator ansatz (QAOA) [20] can be used as a subroutine to build an approximate solver for #P problems. If QAOA succeeds in sampling solutions sufficiently close to uniformly with a finite success rate, then only a polynomial number of samples is needed to obtain an approximate count through the JVV algorithm. Our approach then offers an exponential improvement over previous related work on quantum counting [17]. The key question then becomes: under which conditions can QAOA function as an efficient random sampler of solutions?
@@
This paper is organized as follows. Sec. 2 introduces the quantum alternating operator ansatz, with an emphasis on the quantum approximate optimization algorithm and its Grover-mixer variant, in the context of approximate counting. The JVV algorithm is summarized in Sec. 3. With this theory, Sec. 4 defines the VQCount algorithm. Sec. 5 presents numerical evidence of the performance of VQCount. We first benchmark the performance of VQCount on ##\##NAE3SAT and then push simulations to a higher number of qubits for the ##\##1-in-3SAT problem.
2 Quantum alternating operator ansatz
@@
The product $U_{D}U_{P}$ is applied $p$ times to an initial state |$\ket{\psi_{0}}$, prepared to be an eigenstate of $H_{D}$. For example, |$\ket{\psi_{0}}=\ket{+}^{\otimes n}$ for the drive (1). This yields
@@
In the $p\to\infty$ limit, the convergence of $|\psi(\vec{\beta},\vec{\gamma})\rangle$ to the exact solution of the problem encoded by $H_{P}$ is guaranteed, since $(\vec{\beta},\vec{\gamma})$ can be predetermined to model trotterized adiabatic quantum optimization (AQO) [21]. Clearly, this limit is not interesting for practical use.
@@
where $U_{S}$ is a state preparation unitary and |$\ket{\psi_{0}}=U_{S}\ket{0}^{\otimes n}$. The purpose of $U_{S}$ is to restrict QAOA within a “feasible” subspace of the Hilbert space by eliminating subspaces that are known a priori not to contain solutions.
@@
3 Jerrum-Valiant-Vazirani algorithm
@@
$||P-Q||_{TV}\equiv\frac{1}{2}\sum_{x\in\mathcal{X}}|P(x)-Q(x)|\,.$ (6)
@@
The key element in the proof of JVV is showing that a sample set of polynomial size is sufficient to obtain a multiplicative approximation to the exact solution count.
@@
(Adapted from Theorem 6.4 in [19]) If a self-reducible #P problem on $n$ Boolean variables admits a solution generator with nonuniformity $\eta$, then the solution count can be approximated to within a multiplicative factor $O(\eta n)$ with high probability with only a polynomial number of calls to the generator.
@@
By requiring additionally a fully polynomial almost uniform generator, that is, a generator with $\eta=O(n^{-1})$ whose runtime is polynomial and whose uniformity can be improved by allowing polynomial increases in runtime, JVV provide a fully polynomial-time randomized approximation scheme (FPRAS) for approximate solution counting of self-reducible formulas. That is, for a self-reducible formula over $n$ variables, the JVV algorithm returns, in time polynomial in $n$ and $\varepsilon^{-1}$, a rational number $C\in\mathbb{Q}$ such that
@@
with confidence $41-\delta=3/$ 4. The algorithm requires $O(n^{2}/\varepsilon^{2})$ calls to the generator. Using a bootstrapping method, the failure probability can be reduced arbitrarily at a runtime cost of $\log(1/\delta)$.
4 Variational quantum counting
The main innovation we introduce in this work is the use of the output state |$\ket{\psi(\vec{\beta},\vec{\gamma})}$ of a QAOA circuit, as expressed in Eq. (2), as the generator component of the JVV self-reduction. The result is a variational quantum algorithm for counting.
The discussion in this section will be restricted to the GM-QAOA variant, which, as mentioned in Sec. 2, guarantees that solution amplitudes in |$\ket{\psi(\vec{\beta},\vec{\gamma})}$ are exactly equal. Thus, whenever the outcome of a measurement of the qubits in the computational basis is a solution, then it is sampled uniformly at random from the solution set. In terms of the nonuniformity defined in Eq. 7, we therefore have $\eta=0$. The critical figure of merit for the efficacy of sampling from |$\ket{\psi(\vec{\beta},\vec{\gamma})}$ is hence the success rate of GM-QAOA, that is, the probability that the measurement outcome is a solution, which we can define as
$r=\matrixelement{\psi(\vec{\beta},\vec{\gamma})}{\hat{\cal P}_{\cal G}}{\psi(\vec{\beta},\vec{\gamma})}\,,$ (14)
@@ 4 Variational quantum counting
$\hat{\cal P}_{\cal G}=\sum_{x\in\cal G}\outerproduct{x}{x}\,.$ | . (15)
With this definition, the performance of different sampling-based algorithms for counting can be parameterized on equal footing using the sampling success rate $r$, the multiplicative error $\varepsilon$, and the confidence $1-\delta$. This characterization extends to classical algorithms employing a generator that fails to produce a solution at a rate of $1-r$.
The output state |$\ket{\psi(\vec{\beta},\vec{\gamma})}$ of an optimized GM-QAOA circuit, combined with postselection to eliminate non-solutions, can be used straightforwardly as a solution generator to a self-reducible problem. However, the JVV algorithm requires also a solver for each subproblem. Direct use of the GM-QAOA circuit optimized for the full problem would require additional postselection on the variable fixed in the subproblem.
@@
Guided by the intuition outlined above, we introduce a variational quantum algorithm for approximate counting called VQCount. Its steps are specified as Algorithm 1 below. Given a self-reducible SAT formula $\varphi(x)$, we define the problem Hamiltonian $H_{P}$ whose ground states represent the solutions to $\varphi(x)$. In this work, we map $\varphi(x)$ to the Ising model using standard techniques [30] (see Sec. 5 for details). A GM-QAOA circuit is optimized to minimize the expectation value of $H_{P}$. We then sample solutions by measuring the output of the circuit and postselecting solutions only. By virtue of the GM-QAOA algorithm, solutions are sampled uniformly. Once a sufficient number of samples is collected, we evaluate the conditional probabilities for the two values of a given variable and choose the value corresponding to the largest probability, as in the JVV algorithm. This choice leads to a reduced GM-QAOA circuit, as in Fig. 2, and the process restarts with this circuit (but forgoing parameter optimization), until all variables are fixed.
@@
2:  $\text{PQC}(\vec{\beta},\vec{\gamma})\leftarrow\text{Optimize}(\text{PQC}(\vec{\beta}_{0},\vec{\gamma}_{0}),n_{o})$
3:  $\tilde{N}\leftarrow 1$
@@
Since the foundation of VQCount is the JVV algorithm (as described in Sec. 3), its usage requires $O\left(\frac{n^{2}\log(1/\delta)}{r\varepsilon^{2}}\right)$ samples to approximate the ground-state count to multiplicative accuracy $\varepsilon$ with confidence $1-\delta$, given success rate $r$ of the QAOA circuit, as defined in Eq. (14). This compares favorably to related work on variational algorithms for evaluating the partition function of classical spin Hamiltonians [17], which would require $O\left(\frac{\sqrt{N\log(1/\delta)}}{r\varepsilon}\right)$ samples for the same task. When $N=O(2^{n})$, VQCount requires exponentially fewer samples than the methods of Ref. [17].
A few comments are in order. First, generating quantum states that correspond to arbitrary probability distributions, including uniform distributions over combinatorial structures, is hard in general, and easy only in exceptional cases [31, 32]. Furthermore, in general, GM-QAOA requires circuit depths scaling exponentially with problem size to achieve finite success rates [33]. These observations simply confirm that there is no “free lunch” for quantum counting algorithms: #P-hard problems remain hard for quantum computers, assuming complexity theory hypotheses widely believed to be true. In light of this, one may ask instead whether hybrid algorithms perform well on average in practice. In the next section, we tackle this question for VQCount.
5 Solving #P-Complete Problems with VQCount
@@
Many random SAT formula families exhibit a phase transition phenomenon as a function of the density of clauses $\alpha=m/n$, where $m$ and $n$ denote the numbers of clauses and variables, respectively. The hardness of random #NAE3SAT is maximal close to the satisfiability threshold, located at a critical clause-to-variable ratio $\alpha_{c}\approx 2.1$ [37]. To study the effect of average formula hardness on VQCount performance, we generate instances at $\alpha=1$ and $\alpha=2$. We also restrict the sampled instances to connected biregular graphs. Similarly, positive ##\##1-in-3SAT instances are hardest close to the satisfiability threshold $\alpha_{c}\approx 2/3$ [38]. To study this problem in the hard regime, we generate random cubic graphs, placing a clause on each vertex and a variable on each edge. ##\##1-in-3SAT close to the threshold belongs to the category of locked problems [39], in which the Hamming distance between any two solutions is $O(n)$. #P problems in this regime are arguably the most challenging.
@@
Figure 4: VQCount performance with depth $p=3$ QAOA (blue circles) and GM-QAOA (red squares) circuits for NAE3SAT instances at $\alpha=1$ (left panels) and $\alpha=2$ (right panels). Solid and dotted lines represent the mean and median, respectively. Shaded regions show the standard error of the mean. (a,b) Maximal nonuniformity throughout the self-reduction. The JVV algorithm requires the nonuniformity to be at most $O(n^{-1})$ for $n$ variables, as illustrated (dotted-dashed black lines) with the function $f(n)=O(1/n)$. (c,d) Minimal success rate throughout the self-reduction. (e,f) Numbers of samples needed to achieve an approximate count with error tolerance $\varepsilon=1/3$. An exponential fit extrapolated from the last five points of the median gives $O(1.45^{n})$ (QAOA) and $O(1.44^{n})$ (GM-QAOA) for (e), and $O(1.34^{n})$ (QAOA) and $O(1.88^{n})$ (GM-QAOA) for (f).
@@
The performance of QAOA methods depends sensitively on the choice of $(\vec{\beta}_{0},\vec{\gamma}_{0})$. As the training of VQAs is NP-hard in general [47], a good choice of initial parameters is necessary to avoid optimization pitfalls, such as rugged optimization landscapes and barren plateaus. One method of initialization is trotterized quantum annealing (TQA) [2], which has been found to greatly improve QAOA performance compared to random initialization. We will be using this technique in our simulations.
@@
Figure 5: Number of post-selected samples needed to achieve an approximate count with error tolerance $\varepsilon=1/3$ with depth $p=3$ QAOA (blue circles) and GM-QAOA (red squares) circuits for NAE3SAT instances at $\alpha=1$. Solid and dotted lines represent the mean and median, respectively. Shaded regions show the standard error of the mean. A polynomial fit of the median gives $O(n^{5.12})$ (QAOA) and $O(n^{3.42})$ (GM-QAOA). The inset panel shows the scaling behavior with the inverse of the error tolerance $\varepsilon$ for $n=12$ variables. An polynomial fit of the mean gives $O(\varepsilon^{-1.13})$ (QAOA) and $O(\varepsilon^{-1.07})$ (GM-QAOA).
Figs. 4(e) and 4(f) show the total number of samples needed for VQCount to yield an approximate count within a multiplicative error of $\varepsilon=1/3$. The procedure we use to determine this number is described in Appendix A. The number of samples grows exponentially with instance size due to the postselection of solutions for both QAOA and GM-QAOA as backends for sampling. Despite the exponentially lower success rate of GM-QAOA, VQCount requires roughly equal numbers of samples with either QAOA or GM-QAOA at $\alpha=1$. In contrast, at $\alpha=2$, VQCount with QAOA requires exponentially fewer samples than with GM-QAOA, despite the nonuniformity in solution sampling with QAOA. This suggests that QAOA may be a more efficient sampling backend for VQCount in cases where solutions are few and far apart. We further investigate this behavior in our discussion of #1-in-3SAT below. The number of post-selected samples required for VQCount to produce an estimate within a given factor $\varepsilon$ of the exact count, presented in Fig. 5, appears polynomial in the number of variables and $1/\varepsilon$ as expected.
@@
In this work, the angle parameters of the QAOA circuit are only optimized once and held constant throughout the self-reduction process, as re-optimization at each step would be prohibitively costly. In Sec. 4 we showed that, in the Grover limit, GM-QAOA success rates are non-decreasing with self-reduction step. We now test whether this is also the case away from the Grover limit for GM-QAOA and for regular QAOA in general. Our results are summarized in Fig. 7. We observe that success rates remain non-decreasing for GM-QAOA. In contrast, in QAOA nonuniformity initially increases before decreasing, as seen in Fig. 7(a,b), with the increase being more pronounced for $\alpha=1$. This effect correlates with a decrease in success rate in the first few self-reductions. The success rate of QAOA is nevertheless always superior to that of GM-QAOA.
@@
Figure 8: Scaling behavior of VQCount with QAOA circuits for 1-in-3SAT instances at $\alpha=2/3$. The number of samples needed to achieve an approximate count using naive rejection sampling and the quantum counting algorithm of Brassard, Høyer, Mosca and Tapp (BHMT) are shown with a black dash-dotted line and a black densely dash-dotted line, respectively, following the functions $f(n)=O(1.82^{n})$ and $f(n)=O(1.35^{n})$ for $n$ variables. Solid lines represent the mean, and shaded regions show the standard error of the mean. Dashed lines indicate the exponential fit extrapolated from the last four points as shown in the legend. (a) Number of samples needed for the approximate count to be within error tolerance $\varepsilon=1/3$ for depths $p$. (b) Number of samples needed for the approximate count to be within error tolerance $\varepsilon$ for a depth $p=3$.
@@
Fig. 8 summarizes the scaling behaviour of VQCount with low-depth QAOA. As seen in Fig. 8(a), VQCount scales exponentially better than naive rejection sampling and its performance improves with circuit depth. Moreover, its performance is comparable to that of the quantum counting algorithm by Brassard, Høyer, Mosca and Tapp (BHMT) [13], which scales as $O(\sqrt{\frac{N}{M}})$, where $M$ is the size of the solution space and $N$ is the total size of the search space. At the depths we can access using reasonable computational resources, VQCount scales worse than state-of-the-art classical exact solvers for this problem [23, 24, 25]. The accuracy of VQCount can be improved by increasing the number of samples taken, as shown in Fig. 8(b). Finally, Fig. 9 shows the number of postselected samples needed to produce an estimate within a given factor $\varepsilon$ of the exact count as a function of number of variables and $1/\varepsilon$. Our results suggest a superpolynomially but subexponentially increasing number of postselected samples (cf. Fig. 5 for #NAE3SAT).

### v2

@@ Abstract
We introduce a variational algorithm based on the quantum alternating operator ansatz (QAOA) for the approximate solution of computationally hard counting problems. Our algorithm, dubbed VQCount, is based on the equivalence between random sampling and approximate counting and employs QAOA as a solution sampler. We first prove that VQCount improves upon previous work by reducing exponentially the number of samples needed to obtain an approximation within an arbitrary small multiplicative factor of the exact count. Using tensor network simulations, we then study the typical performance of VQCount with shallow circuits on synthetic instances of two #P-hard problems, positive #NAE3SAT and positive #1-in-3SAT. We employ the original quantum approximate optimization algorithm version of QAOA, as well as the Grover-mixer variant which guarantees a uniform solution probability distribution. We observe a tradeoff between QAOA success probability and sampling uniformity, which we exploit to achieve an empirical efficiency gain over both naive rejection sampling and Grover-based quantum counting. Our results highlight the potential and limitations of variational algorithms for approximate counting.
I Introduction
@@
These considerations motivate us to turn our attention to counting problems. This class of problems asks how many solutions there are for a given instance of a decision problem. Counting problems lie higher in the computational complexity hierarchy than the optimization problems commonly targeted with near-term quantum algorithms. Counting problems arise across many disciplines, with applications in probabilistic reasoning [45, 48, 3], network reliability [53, 20], statistical physics [34], and artificial intelligence [7]. For decision problems whose solutions are verifiable in time that scales polynomially with the size of the problem instance, that is, in polynomial time, the corresponding task of counting solutions belongs to complexity class #P. The class #P, like NP, contains a subset of hard problems which are unlikely to admit exact polynomial-time solutions. Naturally, the counting counterpart of every NP-hard problem in NP is #P-hard. In the quantum domain, counting was originally formulated as an application of amplitude estimation [12, 55, 2], which uses the Grover iteration [27] as a subroutine. More recently, VQAs based on the quantum alternating operator ansatz (QAOA) [29] for (weighted) counting [52] and enumeration [58] were introduced. These algorithms are parameterized by the success rate of QAOA circuits, i.e., the probability of a measurement yielding a solution.
The improbability of exact polynomial-time algorithms for #P-hard problems has spurred extensive research into practical solvers [9]. Exact counting heuristics employ techniques such as component caching [47], knowledge compilation [37], and tensor networks [36, 18] to outperform brute-force enumeration, but they typically remain limited to instances with at most a few hundred variables. Approximate counting algorithms, which sacrifice exactness for performance, fall into two main categories. Hashing-based methods recursively partition the solution space with random hash functions and query an NP oracle, such as a SAT solver, to test for emptiness [51, 16]. Sampling-based methods, by contrast, rely on generating solutions (almost) uniformly at random. An example is the celebrated algorithm by Jerrum, Valiant, and Vazirani (JVV) [33], which establishes an equivalence between approximate counting and random sampling of solutions. The JVV algorithm approximately solves #P-hard problems that are self-reducible if solutions can be sampled sufficiently uniformly (the term “self-reducible” will be defined in Sec. III).
In this work, we introduce VQCount, an algorithm for approximate counting with VQAs. Concretely, we show how QAOA can be used as a subroutine within the JVV algorithm to build an approximate solver for self-reducible #P problems. If QAOA succeeds in sampling solutions sufficiently close to uniformly with success rate $r$, VQCount requires $O\left(\frac{n^{2}\log(1/\delta)}{r\varepsilon^{2}}\right)$ samples to obtain an approximate count to any desired multiplicative accuracy $\varepsilon$ with confidence $1-\delta$ for a problem instance of $n$ variables. Our approach then offers an exponential improvement over previous related work on approximate counting with VQAs [52]. The key question then becomes: under which conditions can QAOA function as an efficient random sampler of solutions?
@@
This paper is organized as follows. Sec. II introduces the quantum alternating operator ansatz, with an emphasis on the quantum approximate optimization algorithm and its Grover-mixer variant, in the context of approximate counting. The JVV algorithm is summarized in Sec. III. With this theory, Sec. IV defines the VQCount algorithm. Sec. V presents numerical evidence of the performance of VQCount. We first benchmark the performance of VQCount on #\#NAE3SAT and then push simulations to a higher number of qubits for the #\#1-in-3SAT problem.
II Quantum alternating operator ansatz
@@
[…]
@@
In the $p\to\infty$ limit, the convergence of $|\psi(\vec{\beta},\vec{\gamma})\rangle$ to the exact solution of the problem encoded by $H_{P}$ is guaranteed, since $(\vec{\beta},\vec{\gamma})$ can be predetermined to model trotterized adiabatic quantum optimization [22]. Clearly, this limit is not interesting for practical use.
@@
[…]
@@
III Jerrum-Valiant-Vazirani algorithm
@@
$P-Q||_{TV}\equiv\frac{1}{2}\sum_{x\in\mathcal{X}}|P(x)-Q(x)|\,.$ (6)
@@
The key element in the proof of the JVV algorithm is showing that a sample set of polynomial size is sufficient to obtain a multiplicative approximation to the exact solution count.
@@
(Adapted from Theorem 6.4 in [33] and Theorem 13.3 in [41]) If a self-reducible #P problem on $n$ Boolean variables admits a solution generator with nonuniformity $\eta^{\prime}$ such that $\eta^{\prime}\leq\eta=O(\varepsilon n^{-1})$, then the solution count can be approximated, with high probability, to within a multiplicative factor of $1+\varepsilon$ using only a polynomial number of calls to the generator.
@@
By requiring additionally a fully polynomial almost uniform generator, that is, a solution generator whose nonuniformity is at most $\eta$ and whose runtime is polynomial in $n$ and $\log(\eta^{-1})$, JVV provides a fully polynomial-time randomized approximation scheme (FPRAS) for approximate solution counting of self-reducible formulas. That is, for a self-reducible formula over $n$ variables, the JVV algorithm returns, in time polynomial in $n$ and $\varepsilon^{-1}$, a rational number $C\in\mathbb{Q}$ such that
@@
with confidence $1-\delta=3/4$. The algorithm requires $O(n^{2}/\varepsilon^{2})$ calls to the generator [35]. Using a bootstrapping method, the failure probability can be reduced arbitrarily at a runtime cost of $\log(1/\delta)$.
IV Variational quantum counting
[…]
@@ IV Variational quantum counting
$|.\hat{\cal P}_{\cal G}=\sum_{x\in\cal G}\outerproduct{x}{x}\,$. (15)
With this definition, the performance of different sampling-based algorithms for counting can be parameterized on equal footing using the sampling success rate $r$, the multiplicative error $\varepsilon$, and the confidence $1-\delta$. This characterization extends to classical algorithms employing a generator that fails to produce a solution at a rate of 1r1-r.
[…]
@@
Guided by the intuition outlined above, we introduce a variational quantum algorithm for approximate counting called VQCount. Its steps are summarized in Algorithm 1 below. Given a self-reducible SAT formula $\varphi(x)$, we define the problem Hamiltonian $H_{P}$ whose ground states represent the solutions to $\varphi(x)$. In this work, we map $\varphi(x)$ to the Ising model using standard techniques [38] (see Sec. V for details). A GM-QAOA circuit is optimized to minimize the expectation value of $H_{P}$. We then sample solutions by measuring the output of the circuit and postselecting solutions only. By virtue of the GM-QAOA algorithm, solutions are sampled uniformly. Once a sufficient number of samples is collected, we evaluate the conditional probabilities for the two values of a given variable and choose the value corresponding to the largest probability, as in the JVV algorithm. This choice leads to a reduced GM-QAOA circuit, as in Fig. 2, and the process restarts with this circuit (but forgoing parameter optimization), until all variables are fixed.
@@
[…]
@@
Since VQCount is built on the JVV algorithm (as described in Sec. III), its usage requires $O\left(\frac{n^{2}\log(1/\delta)}{r\varepsilon^{2}}\right)$ samples to approximate the ground state count to multiplicative accuracy $\varepsilon$ with confidence $1-\delta$, given success rate $r$ of the GM-QAOA circuit, as defined in Eq. (14). This compares favorably to related work on variational algorithms for counting weighted ground states of classical spin Hamiltonians [52]. When all configurations are assigned equal weight, the weighted counting problem reduces to the unweighted case, which requires $O\left(\frac{\sqrt{N\log(1/\delta)}}{r\varepsilon}\right)$ samples for the same task. Consequently, when $N=O(2^{n})$, VQCount requires exponentially fewer samples than the methods of Ref. [52]. Alternative VQA circuits that implement solution sampling may also be used inside VQCount to obtain heuristics to which the theoretical sample-complexity guarantees derived above may not apply. We will investigate such variants in the next section.
A few comments are in order. First, generating quantum states that correspond to arbitrary probability distributions, including uniform distributions over combinatorial structures, is hard in general, and easy only in exceptional cases [28, 5]. Furthermore, in general, GM-QAOA requires circuit depths scaling exponentially with problem size to achieve finite success rates [13, 56]. These observations simply confirm that there is no “free lunch” for quantum counting algorithms: #P-hard problems remain hard for quantum computers, assuming complexity theory hypotheses widely believed to be true. In light of this, one may ask instead whether hybrid algorithms perform well on average in practice. In the next section, we tackle this question for VQCount.
V Solving #P-Complete Problems with VQCount
@@
Figure 4: VQCount performance with depth $p=3$ QAOA (blue circles) and GM-QAOA (red squares) circuits for NAE3SAT instances at $\alpha=1$ (left panels) and $\alpha=2$ (right panels). Solid and dotted lines represent the mean and median, respectively. Shaded regions show the standard error of the mean. (a,b) Maximal nonuniformity throughout the self-reduction. The JVV algorithm requires the nonuniformity to be at most $O(n^{-1})$ for $n$ variables, as illustrated (dotted-dashed black lines) with the function $f(n)\propto 1/n$. (c,d) Minimal success rate throughout the self-reduction. (e,f) Numbers of samples needed to achieve an approximate count with error tolerance $\varepsilon=1/3$. An exponential fit extrapolated from the last five points of the median indicates an asymptotic scaling of $(1.5\pm 0.2)^{n}$ (QAOA) and $(1.44\pm 0.02)^{n}$ (GM-QAOA) for (e), and $(1.34\pm 0.06)^{n}$ (QAOA) and $(1.88\pm 0.06)^{n}$ (GM-QAOA) for (f).
@@
Many random SAT formula families exhibit a phase transition phenomenon as a function of the density of clauses $\alpha=m/n$, where $m$ and $n$ denote the numbers of clauses and variables, respectively. The hardness of random #NAE3SAT is maximal close to the satisfiability threshold, located at a critical clause-to-variable ratio $\alpha_{c}\approx 2.1$ [4]. To study the effect of average formula hardness on VQCount performance, we generate instances at $\alpha=1$ and $\alpha=2$. We also restrict the sampled instances to connected biregular graphs. Similarly, positive #\#1-in-3SAT instances are hardest close to the satisfiability threshold $\alpha_{c}\approx 2/3$ [44]. To study this problem in the hard regime, we generate random cubic graphs, placing a clause on each vertex and a variable on each edge. #\#1-in-3SAT close to the threshold belongs to the category of locked problems [57], in which the Hamming distance between any two solutions is $O(n)$. #P problems in this regime are arguably the most challenging.
@@
Figure 5: Number of post-selected samples needed to achieve an approximate count with error tolerance $\varepsilon=1/3$ with depth $p=3$ QAOA (blue circles) and GM-QAOA (red squares) circuits for NAE3SAT instances at $\alpha=1$. Solid and dotted lines represent the mean and median, respectively. Shaded regions show the standard error of the mean. A polynomial fit of the median indicates an asymptotic scaling of $n^{3.7\pm 0.2}$ (QAOA) and $n^{3.42\pm 0.07}$ (GM-QAOA). The inset panel shows the scaling behavior with the inverse of the error tolerance $\varepsilon$ for $n=12$ variables. A polynomial fit of the mean indicates an asymptotic scaling of $\varepsilon^{-1.10\pm 0.03}$ (QAOA) and $\varepsilon^{-1.08\pm 0.01}$ (GM-QAOA).
@@
The performance of QAOA methods depends sensitively on the choice of $(\vec{\beta}_{0},\vec{\gamma}_{0})$. As the training of VQAs is NP-hard in general [10], a good choice of initial parameters is necessary to avoid optimization pitfalls, such as rugged optimization landscapes and barren plateaus. One method of initialization is trotterized quantum annealing [46], which has been found to greatly improve QAOA performance compared to random initialization. We will be using this technique in our simulations.
@@
Figs. 4(e) and 4(f) show the total number of samples (including those discarded by postselection) needed for VQCount to yield an approximate count within a multiplicative error of $\varepsilon=1/3$. The procedure we use to determine this number is described in Appendix A. The number of samples grows exponentially with instance size due to the postselection of solutions for both QAOA and GM-QAOA as backends for sampling. Despite the exponentially lower success rate of GM-QAOA, VQCount requires roughly equal numbers of samples with either QAOA or GM-QAOA at $\alpha=1$. In contrast, at $\alpha=2$, VQCount with QAOA requires exponentially fewer samples than with GM-QAOA, despite the nonuniformity in solution sampling with QAOA. This suggests that QAOA may be a more efficient sampling backend for VQCount in cases where solutions are few and far apart. We further investigate this behavior in our discussion of #1-in-3SAT below. The number of post-selected samples (excluding those discarded by postselection) required for VQCount to produce an estimate within a given factor $\varepsilon$ of the exact count, presented in Fig. 5, appears polynomial in the number of variables and $1/\varepsilon$ as expected.
@@
In this work, the angle parameters of the QAOA circuit are only optimized once and held constant throughout the self-reduction process. In Sec. IV we showed that, in the Grover limit, GM-QAOA success rates are non-decreasing with self-reduction step. We now test whether this is also the case away from the Grover limit for GM-QAOA and for regular QAOA in general. Our results are summarized in Fig. 7. We observe that success rates remain non-decreasing for GM-QAOA. In contrast, in QAOA nonuniformity initially increases before decreasing, as seen in Fig. 7(a,b), with the increase being more pronounced for $\alpha=1$. This effect correlates with a decrease in success rate in the first few self-reductions. The success rate of QAOA is nevertheless always superior to that of GM-QAOA for the instances we studied. A secondary reason for our choice not to re-optimize parameters is that re-optimization would compound the considerable computational cost of the tensor network simulations required to produce Fig. 7. A study of parameter re-optimization would be more amenable to VQCount runs on quantum hardware, which we hope to perform in future work.
@@
Figure 8: Scaling behavior of VQCount with QAOA circuits for 1-in-3SAT instances at $\alpha=2/3$. The number of samples needed to achieve an approximate count using naive rejection sampling and the BHMT quantum counting algorithm are shown with a black dash-dotted line and a black densely dash-dotted line, respectively, following the functions $f(n)\propto 1.82^{n}$ and $f(n)\propto 1.35^{n}$ for $n$ variables. Solid lines represent the mean, and shaded regions show the standard error of the mean. Dashed lines indicate the exponential fit extrapolated from the last four points as shown in the legend. (a) Number of samples needed for the approximate count to be within error tolerance $\varepsilon=1/3$ for depths pp. The exponential fit indicates an asymptotic scaling of $(1.34\pm 0.02)^{n}$, $(1.28\pm 0.01)^{n}$, and $(1.25\pm 0.01)^{n}$. (b) Number of samples needed for the approximate count to be within error tolerance $\varepsilon$ for a depth $p=3$. The exponential fit indicates an asymptotic scaling of $(1.26\pm 0.01)^{n}$, $(1.28\pm 0.01)^{n}$, and $(1.31\pm 0.01)^{n}$.
@@
Fig. 8 summarizes the scaling behaviour of VQCount with low-depth QAOA. As seen in Fig. 8(a), VQCount scales exponentially better than naive rejection sampling and its performance improves with circuit depth. Moreover, its performance is comparable to that of the quantum counting algorithm by Brassard, Høyer, Mosca and Tapp (BHMT) [12], which scales as $O(\sqrt{\frac{N}{M}})$, where $M$ is the size of the solution space and $N$ is the total size of the search space. At the depths we can access using reasonable computational resources, VQCount scales worse than state-of-the-art classical exact solvers for this problem [36, 18, 19]. The accuracy of VQCount can be improved by increasing the number of samples taken, as shown in Fig. 8(b). Finally, Fig. 9 shows the number of post-selected samples needed to produce an estimate within a given factor $\varepsilon$ of the exact count as a function of number of variables and $1/\varepsilon$. Our results suggest a superpolynomially but subexponentially increasing number of post-selected samples (cf. Fig. 5 for #NAE3SAT).